# Lattice isomorphic Banach lattices of polynomials

### v1

@@ 3 Regular nuclear polynomials
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\sum_{k=1}^{\infty}\frac{6k}{2^{k}}<\infty.$
Defining $F_{k,j}=\left\{{y\in F_{k}}\,:\,{|y|\leq m\,v_{k,j}\,\text{ for some }m\in\mathbb{N}}\right\}$ for all $k,j\in\mathbb{N}$, we get that $((F_{k,j})_{k,j},v_{k,j})$ is a sequence of principal pairs of $F$.
@@
Let $\varepsilon>0$. As $F$ has the LAP, there exists a finite rank operator $T\colon F\to F$ such that $\||(T-id)|_{F_{k,j}}|(v_{k,j})\|<\varepsilon$ for all $k,j\in\mathbb{N}$. Defining $\Psi\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ by $\Psi(z\otimes y)=z\otimes T(y)$ for all $z\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $y\in F$, we get that $\overline{\Phi}_{n}\circ\Psi(u)=T\circ(\overline{\Phi}_{n}(u))$ for every $u\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$, and hence $\overline{\Phi}_{n}(\Psi(w))=T\circ(\overline{\Phi}_{n}(w))=0$. As $\Psi(w_{k})\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$ for every $k\in\mathbb{N}$, $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$ is closed in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ and $\Psi(w)=\displaystyle\sum_{k=1}^{\infty}\Psi(w_{k})$, we get that $\Psi(w)\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$, which yields that $0=\overline{\Phi}_{n}(\Psi(w))=\Phi_{n}(\Psi(w))$, where $\Phi_{n}$ is the following isomorphism of vector spaces:
$\Phi_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F\to\mathcal{P}_{f}(^{n}E;F),\quad\Phi_{n}(\theta_{n}(x^{*})\otimes y)=(x^{*}(x))^{n}y.$
Now, we may proceed as in [4, p.730] to conclude that $\|w\|_{|\pi|}\leq\varepsilon\sum_{k=1}^{\infty}\sum_{j=1}^{n_{k}}\|u_{k,j}\|_{s,|\pi|}$, and hence $w=0$, a contradiction. ∎
@@
Let $E,F,G$ be Banach lattices such that $G$ is Dedekind complete with the LAP. If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
@@
If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$ and $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$ are lattice isomorphic. Indeed, if $\varphi\colon E^{*}\to F^{*}$ is a lattice isomorphism, then $P(x^{*})=\theta_{n}(\varphi(x^{*}))$ defines a positive $n$-homogeneous polynomial from $E$ into $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$, and so $P^{\otimes}\colon\widehat{\otimes}_{k,s,|\pi|}E^{\ast}\to\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$ is a positive linear operator such that $P^{\otimes}(\theta_{n}(x^{*}))=\theta_{n}(\varphi(x^{*}))$ for every $x^{*}\in E^{*}$. Analogously, there exists a positive operator $S\colon\widehat{\otimes}_{k,s,|\pi|}F^{\ast}\to\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$ such that $S(y^{*})=\theta_{n}(\varphi^{-1}(y^{*}))$ for every $y^{*}\in F^{*}$. Finally, since
@@
holds for every $x^{*}\in E^{*}$, we obtain that $S\circ P^{\otimes}(z)=z$ for every $z\in\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$. The same idea also proves that $P^{\otimes}\circ S(z)=z$ for every $z\in\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$. Thus, $(P^{\otimes})^{-1}=S$, and by [1, Theorem 2.15] $P^{\otimes}$ is a lattice isomorphism. An adaptation of the above argument can be used to prove that $\widehat{\otimes}_{k,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}G$ is lattice isomorphic to $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}\,\widehat{\otimes}_{|\pi|}G$ . Now, by applying Theorem 3.2, we get that $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. ∎
@@ C. Boyd
Centro de Matemática, Computação e Cognição
Universidade Federal do ABC
09.210-580 – Santo André – Brazil.
e-mail: colferaiv@gmail.com

### v2

@@ 3 Regular nuclear polynomials
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\sum_{k=1}^{\infty}\frac{2^{n}\cdot 3k}{2^{k}}<\infty.$
Claim 2: For each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, there are $g_{k,j}^{1},\dots,g_{k,j}^{p_{k,j}}\in G_{k}^{+}$ and $\widetilde{v}_{k,j}^{1},\dots,\widetilde{v}_{k,j}^{p_{k,j}}\in F_{k}^{+}$ such that $|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i},$
$\max_{1\leq j\leq m_{k}}\max_{1\leq i\leq p_{k,j}}\|g_{k,j}^{i}\|=\frac{1}{k}\,\text{ and }\,\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\|\widetilde{v}_{k,j}^{i}\|<\infty.$ (2)
First, for each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, there are $\varphi_{k,j}^{1},\dots,\varphi_{k,j}^{p_{k,j}}\in G_{k}^{+}$ such that $z_{k,j}\leq\sum_{i=1}^{p_{k,j}}\theta_{n}(\varphi_{k,j}^{i})$ and
$\sum_{i=1}^{p_{k,j}}\|\varphi_{k,j}\|^{n}\leq\|z_{k,j}\|_{s,|\pi|}+\frac{1}{m_{k}\displaystyle\max_{1\leq j\leq m_{k}}\|y_{k,j}\|\,2^{k}}.$
Next, we note that, as in Claim 1, we can assume that $\varphi_{k,j}\not=0$, and so we define
$g_{k,j}^{i}=\frac{\varphi_{k,j}^{i}}{k\|\varphi_{k,j}^{i}\|}\quad\text{and}\quad\widetilde{v}_{k,j}^{i}:=(k\|\varphi_{k,j}^{i}\|)^{n}y_{k,j}.$
$0\leq g_{k,j}^{i}\in G_{k}$ and $0\leq\widetilde{v}_{k,1}^{i},\dots,\widetilde{v}_{k,m_{k}}^{i}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\leq\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes{\tilde{v}}^{i}_{k,j},\\ \quad\max_{1\leq j\leq m_{k}}\max_{1\leq i\leq p_{k,j}}\|g_{k,j}^{i}\|=\frac{1}{k},$
and
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\|\widetilde{v}_{k,j}^{i}\|$ $=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}k^{n}\|\varphi_{k,j}^{i}\|^{n}\|y_{k,j}\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k^{n}\left(\|z_{k,j}\|_{s,|\pi|}+\frac{1}{m_{k}\max_{1\leq j\leq m_{k}}\|y_{k,j}\|2^{k}}\right)\|y_{k,j}\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k^{n}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|+\sum_{k=1}^{\infty}\frac{k^{n}}{2^{k}}$
$\leq\sum_{k=1}^{\infty}\frac{2^{n}\cdot 3k^{n}}{2^{k}}+\sum_{k=1}^{\infty}\frac{k^{n}}{2^{k}}<\infty.$
For each $k,j,i$, define
$G_{k,j}^{i}:=\left\{{g\in G_{k}}\,:\,{|g|\leq mg_{k,j}^{i}\,\text{ for some }m\in\mathbb{N}}\right\}$
and
$F_{k,j}=\left\{{y\in F_{k}}\,:\,{|y|\leq m\,v_{k,j}\,\text{ for some }m\in\mathbb{N}}\right\}.$
Ordering $(v_{k,j})$ by
$v_{1,1},\dots,v_{1,m_{1}},v_{2,1}\dots,v_{2,m_{2}},\dots$
and $(g_{k,j}^{i})$ by
$g_{1,1}^{1},\dots,g_{1,1}^{p_{1}},g_{1,2}^{1},\dots,g_{1,2}^{p_{1,2}},\dots,g_{1,m_{1}}^{1},\dots,g_{1,m_{1}}^{p_{1,m_{1}}},\dots$
we get that $(G_{k,j}^{i},g_{k,j}^{i})$ and $((F_{k,j}),v_{k,j})$ are sequences of principal pairs of $E^{*}$ and $F$, respectively. Let $\varepsilon>0$. As $E^{*}$ and $F$ have the LAP, there are finite rank operators $S\colon E^{*}\to E^{*}$ and $T\colon F\to F$ such that
$\||(S-id)_{G_{k,j}^{i}}|(g_{k,j}^{i})\|<\varepsilon\,\text{ and }\,\||(T-id)_{F_{k,j}}|(v_{k,j})\|<\varepsilon$
for all $k,j,i$.
Defining $\Psi\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ by $\Psi(\theta_{n}(x^{*})\otimes y)=\theta_{n}(S(x^{*}))\otimes T(y)$ for all $x^{*}\in E^{\ast}$ and $y\in F$, and extending by linearity and continuity, we get that $\Psi(w_{k})\in\otimes_{n,s}\,S(E^{*})\otimes T(F)$ for every $k\in\mathbb{N}$. To see that $\Psi$ is well-defined we note that if $\sum_{i=1}^{k}\theta_{n}(x_{i}^{*})\otimes y_{i}=\sum_{j=1}^{l}\theta_{n}(z_{j}^{*})\otimes w_{j}$, then
$\sum_{i=1}^{k}\theta_{n}(Sx_{i}^{*})\otimes Ty_{i}$ $=\sum_{i=1}^{k}\varphi_{1}^{\otimes}(\theta_{n}(x_{i}^{*}))\otimes Ty_{i}=\varphi_{1}^{\otimes}\otimes T(\sum_{i=1}^{k}\theta_{n}(x_{i}^{*})\otimes y_{i})$
$=\varphi_{1}^{\otimes}\otimes T(\sum_{j=1}^{l}\theta_{n}(z_{j}^{*})\otimes w_{j})=\sum_{j=1}^{l}\varphi_{1}^{\otimes}(\theta_{n}(z_{j}^{*}))\otimes Tw_{j}$
$=\sum_{j=1}^{l}\theta_{n}(Sz_{j}^{*})\otimes Tw_{j},$
where $\varphi_{1}^{\otimes}$ is the linearization of the regular symmetric $n$-linear map $\varphi_{1}:E^{*}\times\cdots\times E^{*}\to\widehat{\otimes}_{n,s,|\pi|}E^{*}$ given by $\varphi_{1}(x_{1}^{*},\dots,x_{n}^{*})=Sx_{1}^{*}\otimes_{s}\cdots\otimes_{s}Sx_{n}^{*}$.
@@
Moreover, as $\otimes_{n,s}S(E^{*})\otimes T(F)$ is closed in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ and $\Psi(w)=\displaystyle\sum_{k=1}^{\infty}\Psi(w_{k})$, we have that $\Psi(w)\in\otimes_{n,s}S(E^{*})\otimes T(F)$, which yields that $\overline{\Phi}_{n}(\Psi(w))=\Phi_{n}(\Psi(w))$, where $\Phi_{n}$ is the following isomorphism of vector spaces:
$\Phi_{n}\colon\otimes_{n,s}E^{\ast}\otimes F\to\mathcal{P}_{f}(^{n}E;F),\quad\Phi_{n}(\theta_{n}(x^{*})\otimes y)=(x^{*}(x))^{n}y.$
Now, considering $\Gamma:\mathcal{P}_{\cal N}(^{n}E;F)\to\mathcal{P}_{\cal N}(^{n}E;F)$ given by
$\Gamma\left(\sum_{i=1}^{\infty}(x_{i}^{*}(\cdot))^{n}y_{i}\right)=\sum_{i=1}^{\infty}(Sx_{i}^{*}(\cdot))^{n}Ty_{i},$
we have that
$\Phi_{n}(\Psi(\theta_{n}(x^{*})\otimes y))=\Phi_{n}(\theta_{n}(Sx^{*})\otimes Ty)=(Sx^{*}(\cdot))^{n}Ty=\Gamma((x^{*}(\cdot))^{n}y)=\Gamma(\Phi_{n}(\theta_{n}(x^{*})\otimes y))$
holds for all $x^{*}\in E^{*}$ and $y\in F$, hence $\Phi_{n}(\Psi(w_{k}))=\Gamma(\Phi_{n}(w_{k}))$ for every $k\in\mathbb{N}$. Therefore
$\Phi_{n}(\Psi(w))$ $=\overline{\Phi}_{n}(\Psi(w))=\sum_{k=1}^{\infty}\overline{\Phi}_{n}(\Psi(w_{k}))=\sum_{k=1}^{\infty}\Phi_{n}(\Psi(w_{k}))$
$=\sum_{k=1}^{\infty}\Gamma(\Phi_{n}(w_{k}))=\Gamma\left(\sum_{k=1}^{\infty}\Phi_{n}(w_{k})\right)=\Gamma\left(\sum_{k=1}^{n}\overline{\Phi}_{n}(w_{k})\right)=\Gamma(\overline{\Phi}_{n}(w))=0.$
Now, since $\Phi_{n}$ is a linear isomorphism, we obtain that $\Psi(w)=0$. From this, we will prove that $w=0$. First, we notice that, for each $k\in\mathbb{N}$
$|(id-\Psi)(w_{k})|$ $\leq|(id-id\otimes T)(w_{k})|+|(id\otimes T-\Psi)(w_{k})|$
$=|id\otimes(id_{F}-T)(w_{k})|+|(id-\otimes^{n}S)\otimes T(w_{k})|$
$\leq|id\otimes(id_{F}-T)|(|w_{k}|)+|(id-\otimes^{n}S)\otimes T|(|w_{k}|)$
$\leq|id\otimes(id_{F}-T)|\left(\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j}\right)+$
$+|(id-\otimes^{n}S)\otimes T|\left(\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i}\right),$
where in the last inequality we used the constructions obtained in Claim 1 and Claim 2 above. We will divide our proof in two steps now. First:
$|id\otimes(id_{F}-T)|\left(\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j}\right)$ $=\sum_{j=1}^{m_{k}}|id\otimes(id_{F}-T)|(u_{k,j}\otimes v_{k,j})$
$=\sum_{j=1}^{m_{k}}u_{k,j}\otimes|(id-T)_{F_{k,j}}|(v_{k,j})$
Second:
$|(id-\otimes^{n}S)\otimes T|$ $\left(\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i}\right)=\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}|(id-\otimes^{n}S)\otimes T|(\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i})$
$=\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(|(id-S)_{G_{k,j}^{i}}|(g_{k,j}^{i}))\otimes T(\widetilde{v}_{k,j}^{i})$
Combining the above, we obtain
$\|w\|_{|\pi|}$ $=\|w-\Psi(w)\|_{|\pi|}=\|(id-\Psi)(\sum_{k=1}^{\infty}w_{k})\|_{|\pi|}\leq\sum_{k=1}^{\infty}\|(id-\Psi)(w_{k})\|_{|\pi|}$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|\||(id-T)_{F_{k,j}}|(v_{k,j})\|+\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\||(id-S)_{G_{k,j}^{i}}|(g_{k,j}^{i})\|^{n}\|T(\widetilde{v}_{k,j}^{i})\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|\varepsilon+\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\varepsilon^{n}\|T\|\|\widetilde{v}_{k,j}^{i}\|.$
Taking $\varepsilon\longrightarrow 0$, we obtain $\|w\|_{|\pi|}=0$, and we have a contradiction. ∎
@@
Let $E,F,G$ be Banach lattices such that $G$ is Dedekind complete and $E^{*},F^{*}$ and $G$ have the LAP. If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic.
@@
If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$ are lattice isomorphic. Indeed, if $\varphi\colon E^{*}\to F^{*}$ is a lattice isomorphism, then $P(x^{*})=\theta_{n}(\varphi(x^{*}))$ defines a positive $n$-homogeneous polynomial from $E$ into $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$, and so $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\to\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$ is a positive linear operator such that $P^{\otimes}(\theta_{n}(x^{*}))=\theta_{n}(\varphi(x^{*}))$ for every $x^{*}\in E^{*}$. Analogously, there exists a positive operator $S\colon\widehat{\otimes}_{n,s,|\pi|}F^{\ast}\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ such that $S(y^{*})=\theta_{n}(\varphi^{-1}(y^{*}))$ for every $y^{*}\in F^{*}$. Finally, since
@@
holds for every $x^{*}\in E^{*}$, we obtain that $S\circ P^{\otimes}(z)=z$ for every $z\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$. The same idea also proves that $P^{\otimes}\circ S(z)=z$ for every $z\in\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$. Thus, $(P^{\otimes})^{-1}=S$, and by [1, Theorem 2.15] $P^{\otimes}$ is a lattice isomorphism. An adaptation of the above argument can be used to prove that $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}G$ is lattice isomorphic to $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}\,\widehat{\otimes}_{|\pi|}G$ . Now, by applying Theorem 3.3, we get that $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. ∎
We conclude the article by providing an example where Corollary 3.4 can be applied:
Example 3.5.
It follows from [12, Remark 4.3] that every atomic Banach lattice with an order continuous norm has the LAP. So, letting $E$ and $F$ be $c_{0},c$ or the AM-space $X$ from Example 2.6, we obtain that $E^{*}$ and $F^{*}$ are both lattice isomorphic to $\ell_{1}$, hence $E^{*}$ and $F^{*}$ are lattice isomorphic with the LAP. Now, if $G$ is an arbitrary atomic Banach lattice with order continuous norm, we conclude by Corollary 3.4 that $\mathcal{P}^{r}_{\cal N}(^{n}E;G)$ and $\mathcal{P}^{r}_{\cal N}(^{n}F;G)$ are lattice isomorphic.
@@ C. Boyd
Departamento de Matemática
Instituto de Ciências Matemáticas e de Computação
Universidade de São Paulo
13566-590 – São Carlos - SP – Brazil
e-mail: viniciusmiranda@icmc.usp.br